# Quantum entanglement correlations in double quark PDFs

### v1

@@ III.1 Corrections to the density matrix due to one gluon emission
×$\hskip 10.00002pt\times\left[\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}-\frac{1}{2}\frac{1}{n^{2}+\Lambda^{2}}-\frac{1}{2}\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\,\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta,\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}\ ,$ (41)
@@ IV Summary
In sec. II we have computed the density matrix over the two-quark LC momentum fractions in an effective LC constituent quark model of the proton intended for not too small $x_{i}\sim 0.1$ or greater, and a low resolution scale $Q_{0}^{2}$. Upon purging the density matrix of its entanglement negativity and the associated quantum correlations, we observe strong modifications of the dPDF in some range of momentum fractions, specifically for asymmetric $x_{1},x_{2}$ far below the bound $x_{1}+x_{2}=1$. In fact, in this regime we find that purging the entanglement negativity turns a suppression of the quantum correlated dPDF (relative to a product of two single quark PDFs) into a strong enhancement. On the other hand, for large $x_{1}+x_{2}$ or $x_{1}\simeq x_{2}$ the effect of such quantum correlations in this model is found to be small whereas classical correlations enforce $f_{qq}(x_{1},x_{2})/f_{q}(x_{1})f_{q}(x_{2})\to 0$ as $x_{1}+x_{2}\to 1$.
In sec. III we derived the corrections to $\rho_{x_{1}x_{2},x_{1}^{\prime}x_{2}^{\prime}}$ due to the emission of one collinear gluon. This corresponds to one step of QCD scale evolution to $Q^{2}>Q_{0}^{2}$ of the entire density matrix, assuming $\alpha_{s}\log Q^{2}/Q_{0}^{2}\ll 1$; on the diagonal we recover the dPDF DGLAP evolution in terms of convolutions of splitting functions and dPDFs. From the density matrix at the new scale $Q^{2}$ we obtain first qualitative predictions for the QCD scale evolution of quantum entanglement in double quark states in the proton. Interestingly, at the higher scale we do observe significant effects due to quantum correlations for nearly symmetric momentum fractions, $x_{1}\simeq x_{2}$, for fairly small $x_{1}+x_{2}\lesssim 0.3$, where their effect is to suppress the dPDF. These initial explorations motivate the development of the complete QCD evolution equations of the full two-quark density matrix, to account for the emission of an entire ladder of partons. This work is in progress.

### v2

@@ III.1 Corrections to the density matrix due to one gluon emission
$\hskip 10.00002pt\times\left[\frac{\vec{n}\cdot\vec{n}^{\prime}}{(n^{2}+\Delta^{2})(n^{\prime 2}+\Delta^{\prime 2})}-\frac{1}{2}\frac{1}{n^{2}+\Lambda^{2}}-\frac{1}{2}\frac{1}{n^{\prime 2}+\Lambda^{\prime 2}}\right]\,\frac{\Psi^{*}_{\text{qqq}}(\xi^{\prime}\eta^{\prime},\eta^{\prime}(1-\xi^{\prime}),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta^{\prime}(1-\eta^{\prime})\xi^{\prime}(1-\xi^{\prime})}}\,\frac{\Psi_{\text{qqq}}(\xi\eta,\eta(1-\xi),\vec{k}_{1},\vec{k}_{2})}{\sqrt{4\eta(1-\eta)\xi(1-\xi)}}\ ,$ (41)
@@ IV Summary
In sec. II we have computed the density matrix over the two-quark LC momentum fractions in an effective LC constituent quark model of the proton intended for not too small $x_{i}\sim 0.1$ or greater, and a low resolution scale $Q_{0}^{2}$. (For an improved description of the proton at low scale one should account for non-perturbative contributions from higher Fock states such as $|qqqg\rangle$ and $|qqqq\bar{q}\rangle$ in the future.) Upon purging the density matrix of its entanglement negativity and the associated quantum correlations, we observe strong modifications of the dPDF in some range of momentum fractions, specifically for asymmetric $x_{1},x_{2}$ far below the bound $x_{1}+x_{2}=1$. In fact, in this regime we find that purging the entanglement negativity turns a suppression of the quantum correlated dPDF (relative to a product of two single quark PDFs) into a strong enhancement. On the other hand, for large $x_{1}+x_{2}$ or $x_{1}\simeq x_{2}$ the effect of such quantum correlations in this model is found to be small whereas classical correlations enforce $f_{qq}(x_{1},x_{2})/f_{q}(x_{1})f_{q}(x_{2})\to 0$ as $x_{1}+x_{2}\to 1$.
In sec. III we derived the corrections to $\rho_{x_{1}x_{2},x_{1}^{\prime}x_{2}^{\prime}}$ due to the emission of one collinear gluon. This corresponds to one step of QCD scale evolution to $Q^{2}>Q_{0}^{2}$ of the entire density matrix, assuming $\alpha_{s}\log Q^{2}/Q_{0}^{2}\ll 1$; on the diagonal we recover the dPDF DGLAP evolution in terms of convolutions of splitting functions and dPDFs. From the density matrix at the new scale $Q^{2}$ we obtain first qualitative predictions for the QCD scale evolution of quantum entanglement in double quark states in the proton. These results are qualitative because we do not resum multiple evolution steps and are therefore limited to unrealistically weak coupling, $\alpha_{s}=0.1$, and large cutoff on the LC momentum of the gluon, $x=0.1$. Even so, interestingly, at the higher scale we do observe significant effects due to quantum correlations for nearly symmetric momentum fractions, $x_{1}\simeq x_{2}$, for fairly small $x_{1}+x_{2}\lesssim 0.3$, where their effect is to suppress the dPDF. These initial explorations motivate the development of the complete QCD evolution equations of the full two-quark density matrix, to account for the emission of an entire ladder of partons. This work is in progress.
V Acknowledgements
We acknowledge support by the DOE Office of Nuclear Physics through Grant DE-SC0002307.